OpenStudy (shuckle):

If 0^2/3 equals 0, why is 0^-2/3 an undefined?

Nnesha (nnesha):

there is a negative sign so you should flip the fraction to convert negative to positive exponent \[\huge\rm x^{-m}=\frac{ 1 }{ x^{m}}\] and use the fact that u can't divide by 0

OpenStudy (shuckle):

But why is it undefined?

Nnesha (nnesha):

you can't divide by if the denominator is 0 then answer would be undefined \[\huge\rm \frac{ 1 }{ 0 }=undefined \]\[\frac{ 0 }{ 1 }=0\]

OpenStudy (shuckle):

Oh. Okay. Thanks. ^-^

Nnesha (nnesha):

np :=)